Data were collected on the top 1,000 financial advisers. Company A had 239 people on the list and another company, Company B, had 121 people on the list. A sample of 16 of the advisers from Company A and 10 of the advisers from Company B showed that the advisers managed many very large accounts with a large variance in the total amount of funds managed. The standard deviation of the amount managed by advisers from Company A was s1 -$581 million. The standard deviation of the amount managed by advisers from Company B was s2 $485 million. Conduct a hypothesis test at α = 0.10 to determine if there is a significant difference in the population variances for the amounts managed by the two companies, what is your conclusion about the variability in the amount of funds managed by advisers from the two firms? State the null and alternative hypotheses. Find the value of the test statistic. (Round your answer to two decimal places.) 1.44 Find the p-value. (Round your answer to four decimal places.) p-value0.1623Data we
Answer #1

Solution :

This is the two tailed test .

Test statistic = 1.44

Using F table, P-value is

DF of A = 16 - 1 = 15

DF of B = 10 - 1 = 9

P-value = 0.2953